The chord of contact of the pair of tangents drawn from each point on the line 2x+3y=4 to the circle x^2+y^2 =1 pass through which point

Grade:11

2 Answers

ayush
7 years ago
Let any point on line be ((4-3t)/2,tThus equation of chord of contact given by: [T=0] will be [X(4-3t)/2 + Yt - 1=0] which will always pass through (1/2,3/4).
mycroft holmes
7 years ago
You are looking for the pole of the line 2x+3y=4 w.r.t to the circle x2+y2 = 1.
 
If the point is (x1, y1), then the polar of the point would be xx1+yy1 = 1. Comparing with 2x+3y = 4, we immediately get 
\frac{x_1}{2} = \frac{y_1}{3} =\frac{1}{4}
 
so that x1 = ½, and y1 = ¾